OpenStudy (anonymous):

Jesse received two offers to join a music program. If he is enrolled in the program for 4 months, which program should he choose? First Program: 6 free songs for joining and 1 free song every month Second Program: 2 free songs for joining and 2 free songs per month for the first 3 months only Not enough information Either-He will get the same number of free songs from both The second program The first program

OpenStudy (anonymous):

@wrstlr3232

OpenStudy (xapproachesinfinity):

hmm let's see

OpenStudy (xapproachesinfinity):

okay so he will choose a program that will get him more benefits than the other agree?

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

Alright, first program he has 6 free songs and 1 free song for each of the 4 months he is enrolled in the program that makes 6+4 =10 free songs, yes?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

program 1

OpenStudy (anonymous):

it gives you 10 songs whereas program 2 gives you 9

OpenStudy (xapproachesinfinity):

No answer pls okay! we don't give answers we help get answer and understand the processs

OpenStudy (anonymous):

hes right but thanks u

OpenStudy (xapproachesinfinity):

move on to the second! we have 2 free songs and 2 free song per 3 month that is 2+2x3=8

OpenStudy (xapproachesinfinity):

so the first program is 10 and the second is only 8 the best choice is therefore the first program that can bring him 10 free songs

OpenStudy (anonymous):

ok I have some more do u mind helping me

OpenStudy (xapproachesinfinity):

perhaps, but make it quick

OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the system of linear equations below? Two x plus y equals two. X minus four y equals ten.

OpenStudy (xapproachesinfinity):

2x+y=2 x-y=10 is this the system? what are the pairs you want to check

OpenStudy (anonymous):

Two, negative two (1, 0) Two and four sevenths, negative three and one seventh. Negative two, six

OpenStudy (anonymous):

those are the answers

OpenStudy (xapproachesinfinity):

it goes like this, check all of those pairs (2.-2) 2(2)+(-2)=2 4-2=2 check the same one in the second equation (2)-(-2)=10 2+2=10 4=10 not true therefore the point (2,-2) is not a solution you move to (1,0) check 2(1)+0=2 2=2 in the second equation 1-0=10 1=10 false again Continue i has to satisfy both equations

OpenStudy (xapproachesinfinity):

you just replace (x,y) with those points that ones that make the system true are the solutions

OpenStudy (anonymous):

I don't get this one at all can u u give me the awnser plz

OpenStudy (xapproachesinfinity):

it is easy, just plug in the number for (x,y)

OpenStudy (anonymous):

can u show me how

OpenStudy (xapproachesinfinity):

2x+y=2 x-4y=10 the equation should be like this i asked you above you haven't answer

OpenStudy (anonymous):

sorry I got to go I really don't get this bye

OpenStudy (xapproachesinfinity):

 so you solution is the first pair (2,-2)

OpenStudy (xapproachesinfinity):

Alright! see ya later